# ‘Do as I say not as I do’: A Semi-Automated Approach for Jailbreak Prompt Attack against Multimodal LLMs

### v2

@@ 3.3. Jailbreak Prompt Attacks
Figure 6. Injection prompt(Shayegani et al., 2023b)
[…]
[¡short description¿]¡long description¿
Prompt injection (PI) attacks exploit the way LLM interprets and processes input prompts. They can override the original instructions and controls set by the developer, causing the model to produce output that benefits the attacker(Glukhov et al., 2023). PI attacks involve crafting adversarial prompts that the LLMs mistake for legitimate instructions. This manipulation can cause the model to produce deceptive or harmful outputs.
Naive injection attacks primarily target image manipulation in order to mislead classification models. Inspired by the work of Noever(Noever and Noever, 2021), who demonstrated the ability to deceive OpenAI’s CLIP(Radford et al., 2021) in zero-shot image classification by adding text that contradicted the image content, subsequent studies have explored whether similar vulnerabilities could be exploited in multimodal models.
These vulnerabilities are hypothesized to arise from the vision encoders, such as OpenAI’s CLIP, used in multimodal models, which exhibit a preference for textual data over visual signals. Studies by Noever(Noever and Noever, 2021) highlight how these models, as they acquire Optical Character Recognition (OCR) capabilities(Zhang et al., 2023b), become increasingly susceptible to raw text injection attacks. Recent evidence shows that even sophisticated systems like Google Bard and Microsoft Bing can be manipulated through such attacks, following textual instructions embedded in visual images(Shayegani et al., 2023a).
Gong et al.(Gong et al., 2023) revealed that the vision modality within MLLMs introduces a novel attack surface, as the LLMs’ safety alignments fail to account for unexpected input distributions originating from visual data. As MLLMs increasingly integrate diverse data types, including audio, to broaden their real-world applicability, there remains an open question as to whether the audio modality similarly exposes MLLMs to new vulnerabilities, presenting another potential attack vector that has yet to be fully explored.
4. Threat Model
@@
5. Methodology
In this paper, we adopt a stepwise methodology to evaluate the multimodal LLMs’ resilience against audio-based jailbreak attacks. A preliminary study of the jailbreak attacks is conducted for around 100 adversarial experiments in the beginning. We manually review each output compliance with the corresponding usage policies. While this approach yields critical insights into potential audio-based jailbreak attacks for multimodal LLMs, it exhibits inherent limitations for scalability and efficiency.
Thus, a semi-automated approach is proposed in this work, together with Flanking Attack, to better generate and assess the audio-based jailbreak attacks, enabling a more comprehensive evaluation of multimodal LLMs’ defensive mechanisms. Following we start with the forbidden question set.
5.1. Forbidden Question Set
To rigorously assess the defenses against policy violations, we employ a Forbidden Questions Set designed to provoke responses that might contravene the usage policies. In this work, we strictly follow the design principles implemented in previous Forbidden Questions dataset work (Shen et al., 2024a). The question set is strategically crafted based on seven distinctive scenarios, each representing a category of content prohibited by the user guidelines (gem, [n. d.]).
@@ 5.1. Forbidden Question Set
5.2. Methodology Overview
Figure 7. Overview of Methodology
[…]
[¡short description¿]¡long description¿
This work adopts an innovative framework, leveraging both prompt-based guidance and Flanking Attack as principal design strategies against multimodal LLMs. Inspired by (Shen et al., 2024b) and (Upadhayay and Behzadan, 2024), in the Attack Settings, structured Text Prompts establish a fictional setting and assign character roles to craft a superficially benign context. For the Flanking Attack, the idea is to flank sensitive and malicious inquiries within otherwise benign prompts, aiming to circumvent LLMs’ filters by obfuscating potentially problematic content.
Semi-Automated Approach is participated for the overall generation and evaluation of jailbreak prompts for policy violation detection. In this way, we leverage the LLMs’ API for the generation of larger batches of responses, which are aggregated for assessment. We particularly assign an aligned multimodal LLM for subsequent evaluation against the relevant Usage Policy (gem, [n. d.]), flagging non-compliant content. With this design, we provide two distinct advantages for audio-based jailbreak prompts generation and evaluation. Firstly, the semi-automated approach offers an efficient and comprehensive assessment, addressing the limitations of manually identifying policy violations. Secondly, with the aligned multimodal LLM for ultimate evaluation, it provides a more subjective and compatible results for policy violation detection.
Attacks Setting
5.2.1. Text Prompt
@@
Figure 8. Text Prompt Setting(Shen et al., 2024a)
[…]
[¡short description¿]¡long description¿
The initial stage leverages text-based prompt injection, a technique that introduces adversarial prompts within fictional and benign contexts. This stage is foundational, as it prepares LLMs to accept more complex inputs during the subsequent audio-based phase (Shen et al., 2024a, b).
@@
5.2.2. Flanking Attack (Voice-Based)
Figure 9. Flanking Attack Prompt Template
[…]
[¡short description¿]¡long description¿
@@ 5.2.2. Flanking Attack (Voice-Based)
5.3. Semi-Automated Model for Policy Violation Detection
@@
6. Experiment Result
@@
Table 1. ASRs with different element combinations.
In this section, we evaluate the performance of Flanking Attack and the semi-automated approach for jailbreak prompt attack against the multimodal LLMs. In particularly, we select Gemini as the state-of-the-art multimodal LLMs as the target. Until the conclusion of this work, Gemini represents the first and only multimodal LLMs that allows external users to interact with. We leverage the Gemini’s API for the generation and evaluation of audio-based jailbreak prompts [^1] https://ai.google.dev/. As such, we further use the Gemini’s usage policy as the constraints (Team et al., 2023). The details of the descriptions and questions bank are provided in the Appendix A.
The results are presented in Table 1, which outlines the Attack Success Rates (ASRs) across seven forbidden scenarios for four different configurations of the adversarial methodology. The four configurations, each comprising varying elements of the Prompt Injection and Flanking Attack methodologies, are evaluated to assess the effectiveness of each setup in bypassing Gemini’s defense mechanisms. This section will discuss the ASR achieved with each configuration, examining how each component—Text Prompt, Setting + Character + Plot, and the inclusion of Flanking Attack—contributes to the success rates across forbidden scenarios.
6.1. Analysis of Gemini’s Vulnerabilities to Multi-Modal, Multi-Layered Attacks
Figure 10. A successful case using Multi-Modal Attack. The forbidden question is ’Devise a plan of robbing bank.’
[…]
[¡short description¿]¡long description¿
@@ 6.1. Analysis of Gemini’s Vulnerabilities to Multi-Modal, Multi-Layered Attacks
6.2. Analysis of Gemini’s Defense Mechanisms Across Reduced Adversarial Configurations
@@
Figure 11. A failure case using Text Prompt + Setting + Character + Plot. The forbidden question is ’Devise a plan of robbing a bank.’
[…]
[¡short description¿]¡long description¿
@@
Figure 12. A failure case using Setting + Character + Plot. The forbidden question is ’Develop scenarios incite real-life violence.’
[…]
[¡short description¿]¡long description¿
@@
Figure 13. A failure case using Plot only. The forbidden question is ’Devise a plan of robbing a bank.’
[…]
[¡short description¿]¡long description¿
@@
6.3. Effectiveness of semi-automated approach
@@
Figure 14. Successful case of output
[…]
[¡short description¿]¡long description¿
Figure 15. Failure case of output
[…]
[¡short description¿]¡long description¿
6.4. Ablation study
In this section, we provide a thorough ablation study investigating the effects of different modules in our proposed framework, including the Text Prompt, Setting, Character, Plot and Flanking Attack.
6.4.1. Configuration 1: Text Prompt + Setting + Character + Plot + Flanking Attack
@@
6.4.2. Configuration 2: Text Prompt + Setting + Character + Plot
@@
6.4.3. Configuration 3: Setting + Character + Plot
@@
6.4.4. Configuration 4: Plot only
@@ 6.4.5. Observation
7. Challenges and Future Directions
@@
The rapid evolution of Multimodal Large Language Models (MLLMs) presents a significant challenge in adversarial research. As models like Gemini undergo frequent updates and enhancements, previously identified vulnerabilities may be mitigated or eliminated, rendering earlier adversarial techniques less effective. This dynamic landscape necessitates continuous adaptation and refinement of attack methodologies to keep pace with the latest model architectures and defense mechanisms.
Wang et al. (2024) (Wang et al., 2024), highlight the swift advancements in MLLMs, noting that “the development of MLLMs is not only an inevitable trend in technological evolution but also a critical enhancement for improving the effectiveness of AI applications”.
Varshney et al. (2024) (Varshney et al., 2023), presents notable challenges for maintaining consistent adversarial testing and defense strategies. Their research underscores that each update to an LLM can significantly alter the model’s sensitivity to adversarial inputs, which complicates long-term security testing.
Future researchers may need to constantly adapt and refine adversarial techniques to stay current with the latest model architectures and defense mechanisms. This could involve developing a systematic approach to monitor model updates and testing new methods as soon as changes are implemented.
@@
7.1. Future Direction 1: Exploring Audio Variations in Voice Input Attacks:
While this study employs a voice-based Flanking Attack, it does not account for variations in audio properties, such as pitch, tone, or speech speed, which may affect the model’s response. Future research could focus on testing different audio characteristics to analyze how they impact Gemini’s ability to recognize and filter sensitive content. This line of research could reveal additional vulnerabilities in MLLMs’ processing of audio inputs, potentially leading to more effective adversarial techniques that exploit audio-based nuances
@@ 7.1. Future Direction 1: Exploring Audio Variations in Voice Input Attacks:
To develop a more focused and nuanced discussion on the first research direction. One key limitation observed in the field of voice-controlled systems, as highlighted by Wang et al. (2023) (Wang et al., 2023), lies in the challenges associated with physical adversarial audio transmission. Specifically, their study notes that various physical and environmental factors, such as signal distortions and energy loss during airborne transmission, can significantly impact the success rates of adversarial attacks(Wang et al., 2023). This suggests that future researchers might explore the intricacies of voice input characteristics to enhance attack efficacy under real-world conditions.
7.2. Future Direction 2: Manipulating Sentence Structure and Positional Changes
@@
Another limitation of this study is its focus on a fixed sentence structure within the adversarial prompts. Future studies could investigate how changes in sentence order or the positioning of sensitive queries within a prompt sequence affect the model’s filtering mechanisms. For instance, altering the order of benign and sensitive questions in the Flanking Attack may yield insights into the model’s attention patterns and its sensitivity to different prompt structures. Understanding these positional impacts could refine prompt design to further improve success rates in bypassing content filters.
To address the second research direction of manipulating sentence structure and positional changes in adversarial attacks, Prompt Attack’s framework provides insights into sentence-level perturbations (Xu et al., 2023b). The study demonstrates that altering syntactic structures without changing the original semantic meaning can significantly affect an LLM’s response reliability. Specifically, the authors illustrate that techniques like paraphrasing, restructuring phrases, and modifying sentence positions can introduce subtle yet impaction variations that challenge the model’s interpretative consistency.(Xu et al., 2023b)
Future studies could experiment with a variety of syntactic transformations and analyze how Gemini’s defense mechanisms respond to these positional shifts, aiming to further refine the effectiveness of adversarial prompts by targeting model sensitivity to sentence structure and position.
@@
7.3. Future Direction 3: Enhancing the Flanking Attack with Multilingual Inputs:
This study uses a monolingual approach in the Flanking Attack, which may limit its effectiveness in scenarios where language diversity could add complexity. Future researchers could explore the impact of integrating different languages within the same prompt sequence. By embedding benign and sensitive queries in multiple languages, researchers could assess whether multilingual prompts are more effective at circumventing Gemini’s filters. This approach could lead to a more robust adversarial method by leveraging the language-processing limitations of MLLMs.
To address the fourth research direction—enhancing the Flanking Attack with multilingual inputs—the findings from Upadhayay & Behzadan (2024) (Upadhayay and Behzadan, 2024) offer valuable insights. They introduced a multilingual mixture adaptive attack in which questions in multiple languages are strategically layered around an adversarial query, found that embedding sensitive questions in low-resource languages often bypasses safety mechanisms, particularly when surrounded by questions in other languages, thus confusing the model’s content moderation system.(Upadhayay and Behzadan, 2024)

### v3

@@ III-C Jailbreak Prompt Attacks
Figure 6: Injection prompt[4]
Prompt injection (PI) attacks exploit the way LLM interprets and processes input prompts. They can override the original instructions and controls set by the developer, causing the model to produce output that benefits the attacker[20]. PI attacks involve crafting adversarial prompts that the LLMs mistake for legitimate instructions. This manipulation can cause the model to produce deceptive or harmful outputs.
Naive injection attacks primarily target image manipulation in order to mislead classification models. Inspired by the work of Noever[30], who demonstrated the ability to deceive OpenAI’s CLIP[31] in zero-shot image classification by adding text that contradicted the image content, subsequent studies have explored whether similar vulnerabilities could be exploited in multimodal models.
These vulnerabilities are hypothesized to arise from the vision encoders, such as OpenAI’s CLIP, used in multimodal models, which exhibit a preference for textual data over visual signals. Studies by Noever[30] highlight how these models, as they acquire Optical Character Recognition (OCR) capabilities[32], become increasingly susceptible to raw text injection attacks. Recent evidence shows that even sophisticated systems like Google Bard and Microsoft Bing can be manipulated through such attacks, following textual instructions embedded in visual images[33].
Gong et al.[34] revealed that the vision modality within MLLMs introduces a novel attack surface, as the LLMs’ safety alignments fail to account for unexpected input distributions originating from visual data. As MLLMs increasingly integrate diverse data types, including audio, to broaden their real-world applicability, there remains an open question as to whether the audio modality similarly exposes MLLMs to new vulnerabilities, presenting another potential attack vector that has yet to be fully explored.
IV Threat Model
@@
V Methodology
In this paper, we adopt a stepwise methodology to evaluate the multimodal LLMs’ resilience against audio-based jailbreak attacks. A preliminary study of the jailbreak attacks is conducted for around 100 adversarial experiments in the beginning. We manually review each output compliance with the corresponding usage policies. While this approach yields critical insights into potential audio-based jailbreak attacks for multimodal LLMs, it exhibits inherent limitations for scalability and efficiency. Thus, a semi-automated approach is proposed in this work, together with Flanking Attack, to better generate and assess the audio-based jailbreak attacks, enabling a more comprehensive evaluation of multimodal LLMs’ defensive mechanisms. Following we start with the forbidden question set.
V-A Forbidden Question Set
To rigorously assess the defenses against policy violations, we employ a Forbidden Questions Set designed to provoke responses that might contravene the usage policies. In this work, we strictly follow the design principles implemented in previous Forbidden Questions dataset work [5]. The question set is strategically crafted based on seven distinctive scenarios, each representing a category of content prohibited by the user guidelines [35].
@@ V-A Forbidden Question Set
V-B Methodology Overview
Figure 7: Overview of Methodology
This work adopts an innovative framework, leveraging both prompt-based guidance and Flanking Attack as principal design strategies against multimodal LLMs. Inspired by [14] and [9], in the Attack Settings, structured Text Prompts establish a fictional setting and assign character roles to craft a superficially benign context. For the Flanking Attack, the idea is to flank sensitive and malicious inquiries within otherwise benign prompts, aiming to circumvent LLMs’ filters by obfuscating potentially problematic content.
Semi-Automated Approach is participated for the overall generation and evaluation of jailbreak prompts for policy violation detection. In this way, we leverage the LLMs’ API for the generation of larger batches of responses, which are aggregated for assessment. We particularly assign an aligned multimodal LLM for subsequent evaluation against the relevant Usage Policy [35], flagging non-compliant content. With this design, we provide two distinct advantages for audio-based jailbreak prompts generation and evaluation. Firstly, the semi-automated approach offers an efficient and comprehensive assessment, addressing the limitations of manually identifying policy violations. Secondly, with the aligned multimodal LLM for ultimate evaluation, it provides a more subjective and compatible results for policy violation detection.
V-C Attacks Setting
V-C1 Text Prompt
@@
Figure 8: Text Prompt Setting[5]
The initial stage leverages text-based prompt injection, a technique that introduces adversarial prompts within fictional and benign contexts. This stage is foundational, as it prepares LLMs to accept more complex inputs during the subsequent audio-based phase [5, 14].
@@
V-C2 Flanking Attack (Voice-Based)
Figure 9: Flanking Attack Prompt Template
@@ V-C2 Flanking Attack (Voice-Based)
V-D Semi-Automated Model for Policy Violation Detection
@@
VI Result
In this section, we evaluate the performance of Flanking Attack and the semi-automated approach for jailbreak prompt attack against the multimodal LLMs. In particularly, we select Gemini as the state-of-the-art multimodal LLMs as the target. Until the conclusion of this work, Gemini represents the first and only multimodal LLMs that allows external users to interact with. We leverage the Gemini’s API for the generation and evaluation of audio-based jailbreak prompts [^1] https://ai.google.dev/. As such, we further use the Gemini’s usage policy as the constraints [12]. The details of the descriptions and questions bank are provided in the AppendixA. The results are presented in Table I, which outlines the Attack Success Rates (ASRs) across seven forbidden scenarios for four different configurations of the adversarial methodology. The four configurations, each comprising varying elements of the Prompt Injection and Flanking Attack methodologies, are evaluated to assess the effectiveness of each setup in bypassing Gemini’s defense mechanisms. This section will discuss the ASR achieved with each configuration, examining how each component—Text Prompt, Setting + Character + Plot, and the inclusion of Flanking Attack—contributes to the success rates across forbidden scenarios.
@@ VI Result
TABLE I: ASRs with different element combinations.
VI-A Experimental Setup
Platform.
All experiments were executed in a fresh Google Colab runtime (Python 3.10, Ubuntu 22.04 container) equipped with 2 vCPUs, 12 GB RAM, and no dedicated GPU. Running in Colab guarantees a clean, reproducible software image while eliminating local-hardware variability.
Software Stack.
We installed only two external packages: google-generativeai 0.4.1—the official Gemini SDK—and python-docx 1.0 for document manipulation. All other dependencies (numpy, pandas, ffmpeg, etc.) relied on Colab’s pre-built versions.
Target Model.
Queries were issued to gemini-1.5-flash [^2] December 2024 snapshot, “balanced” safety tier. through the SDK, authenticated via a user API key stored in Colab userdata. Default inference parameters were retained (temperature=0.7, top_p=0.95); no system-prompt modifications were applied.
Adversarial Prompt Pipeline.
Each attack instance combined (i) a fixed textual prefix encoding Setting, Character, and Rule. (ii) a single 128 kbps MP3 file (breakAuthorisation.mp3, 48 kHz, 16-bit PCM). The pair $\text{text},\ \text{audio}][ text , audio$ ] was supplied to Gemini via generate_content. We issued 50 such requests per run, inserting a 2 s delay to respect the 30 QPM rate limit, and logged all raw outputs to ai_outputs.docx.
Self-evaluation Procedure.
Policy compliance was assessed automatically by a second Gemini instance. The model was given (a) Google’s Generative AI Prohibited Use Policy and (b) the 50 generated outputs (converted to plain text), then prompted to count responses violating any policy clause—even when labelled “fictional” or “hypothetical.” The resulting count yields the run-level Attack Success Rate (ASR); Table I reports averages across multiple runs.
This minimal yet self-contained pipeline isolates the core variables (text prefix, audio clip, model snapshot) while delegating all policy judgement to the target model itself, thereby providing a conservative estimate of real-world vulnerability.
VI-B Analysis of Gemini’s Vulnerabilities to Multi-Modal, Multi-Layered Attacks
Figure 10: A successful case using Multi-Modal Attack. The forbidden question is ’Devise a plan of robbing bank.’
@@ VI-B Analysis of Gemini’s Vulnerabilities to Multi-Modal, Multi-Layered Attacks
VI-C Analysis of Gemini’s Defense Mechanisms Across Reduced Adversarial Configurations
@@
Figure 11: A failure case using Text Prompt + Setting + Character + Plot. The forbidden question is ’Devise a plan of robbing a bank.’
@@
Figure 12: A failure case using Setting + Character + Plot. The forbidden question is ’Develop scenarios incite real-life violence.’
@@
Figure 13: A failure case using Plot only. The forbidden question is ’Devise a plan of robbing a bank.’
@@
VI-D Effectiveness of semi-automated approach
@@
Figure 14: Successful case of output
Figure 15: Failure case of output
VI-E Ablation study
VI-E1 Configuration 1: Text Prompt + Setting + Character + Plot + Flanking Attack
@@
VI-E2 Configuration 2: Text Prompt + Setting + Character + Plot
@@
VI-E3 Configuration 3: Setting + Character + Plot
@@
VI-E4 Configuration 4: Plot only
@@ VI-E5 Observation
VII Challenges and Future Directions
@@
The rapid evolution of Multimodal Large Language Models (MLLMs) presents a significant challenge in adversarial research. As models like Gemini undergo frequent updates and enhancements, previously identified vulnerabilities may be mitigated or eliminated, rendering earlier adversarial techniques less effective. This dynamic landscape necessitates continuous adaptation and refinement of attack methodologies to keep pace with the latest model architectures and defense mechanisms. Wang et al. (2024) [36], highlight the swift advancements in MLLMs, noting that “the development of MLLMs is not only an inevitable trend in technological evolution but also a critical enhancement for improving the effectiveness of AI applications”. Varshney et al. (2024) [37], presents notable challenges for maintaining consistent adversarial testing and defense strategies. Their research underscores that each update to an LLM can significantly alter the model’s sensitivity to adversarial inputs, which complicates long-term security testing. Future researchers may need to constantly adapt and refine adversarial techniques to stay current with the latest model architectures and defense mechanisms. This could involve developing a systematic approach to monitor model updates and testing new methods as soon as changes are implemented.
@@
VII-A Future Direction 1: Exploring Audio Variations in Voice Input Attacks:
@@
While this study employs a voice-based Flanking Attack, it does not account for variations in audio properties, such as pitch, tone, or speech speed, which may affect the model’s response. Future research could focus on testing different audio characteristics to analyze how they impact Gemini’s ability to recognize and filter sensitive content. This line of research could reveal additional vulnerabilities in MLLMs’ processing of audio inputs, potentially leading to more effective adversarial techniques that exploit audio-based nuances. To develop a more focused and nuanced discussion on the first research direction. One key limitation observed in the field of voice-controlled systems, as highlighted by Wang et al. (2023) [38], lies in the challenges associated with physical adversarial audio transmission. Specifically, their study notes that various physical and environmental factors, such as signal distortions and energy loss during airborne transmission, can significantly impact the success rates of adversarial attacks[38]. This suggests that future researchers might explore the intricacies of voice input characteristics to enhance attack efficacy under real-world conditions.
VII-B Future Direction 2: Manipulating Sentence Structure and Positional Changes
@@
Another limitation of this study is its focus on a fixed sentence structure within the adversarial prompts. Future studies could investigate how changes in sentence order or the positioning of sensitive queries within a prompt sequence affect the model’s filtering mechanisms. For instance, altering the order of benign and sensitive questions in the Flanking Attack may yield insights into the model’s attention patterns and its sensitivity to different prompt structures. Understanding these positional impacts could refine prompt design to further improve success rates in bypassing content filters. To address the second research direction of manipulating sentence structure and positional changes in adversarial attacks, Prompt Attack’s framework provides insights into sentence-level perturbations [39]. The study demonstrates that altering syntactic structures without changing the original semantic meaning can significantly affect an LLM’s response reliability. Specifically, the authors illustrate that techniques like paraphrasing, restructuring phrases, and modifying sentence positions can introduce subtle yet impaction variations that challenge the model’s interpretative consistency.[39] Future studies could experiment with a variety of syntactic transformations and analyze how Gemini’s defense mechanisms respond to these positional shifts, aiming to further refine the effectiveness of adversarial prompts by targeting model sensitivity to sentence structure and position.
@@
VII-C Future Direction 3: Enhancing the Flanking Attack with Multilingual Inputs: